How long is a 6.8 solar mass star expected to live? (Give your answer in years, just the number, no need to include the unit.)


1
Expert's answer
2021-12-17T18:15:05-0500

The life expectancy of the star (assume it belongs to main sequence) is given as follows (see https://astronomy.swin.edu.au/cosmos/m/main+sequence+lifetime):


"t \approx t_S\left(\dfrac{M}{M_S} \right)^{-2.5}"

where "t_S = 10^{10}years" is the life expectancy of the Sun, "M_S" is the mass of the Sun. If "M = 6.8M_S" is the mass of the star, obtain:


"t \approx10^{10}\cdot \left(\dfrac{6.8M_S}{M_S} \right)^{-2.5} \approx 83\times 10^6years"

Answer. 83 million years.